# Mission-Aligned Learning-Informed Control of Autonomous Systems: Formulation and Foundations

### v1

@@ 1 Introduction
In service of this task, the robot is able to lean to roll forward and back as well as turn in full rotation, and rotate its arm along three axes. Assume, for simplicity, that objects can be picked up without modeling grasping, as if everything relevant was magnetized. Objects in the environment define collisions to be avoided, which requires including the speed and position in the state space. The robot has a propulsion system, which is its forward control, and a braking mechanism. The manufacturing of this propulsion system involves some new features which are not system-identified to full certainty, and there is a small but noticeable nonlinear gyroscopic effect on its motion. This effect will have to be learned in the course of its operation.
@@
This work will present a formalism that defines a bilevel optimization problem incorporating Reinforcement Learning (RL)-assisted Model Predictive Control in the bottom layer, and RL-assisted classical planning at the higher level. In describing the formalism we will return to this example, and in ongoing works analyzing the theoretical properties and testing a simulation of this stylized example.
@@
Human-aligned and safe autonomous robotics is a technology of significant interest to many researchers, companies, and governments, with population aging creating many potential labor shortages for assistance care and health for the elderly, the potential for automated manufacturing, and other reasons. As it is also a representative example of the conceptual components of such an operation, it will be a running example throughout this line of work. In this work we simplify some of these concerns and focus on developing the foundations of the mathematical formulation for the optimization framework. We present a formal schema to define the autonomous decision procedures of an agent using multiple relevant disciplines associated with mathematical engineering.
@@
Designing and managing systems towards optimality or looser satisfaction criteria over time is a generic cross-disciplinary problem that has attracted attention from numerous communities. In the case wherein the model of the states and actions can be well represented by a finite set while the system dynamics and the utility of different circumstances is unknown, Reinforcement Learning (RL) has risen to prominence as the preeminent methodology. RL, like Classification (e.g., image recognition), and Generative Learning (e.g., Large Language Models), has been a central component in the accelerating, groundbreaking developments of AI that have characterized the technological landscape for the past decade. Neural networks leverage the raw computational power of highly parallel GPU hardware permitting training of deep models with millions to billions of trainable parameters. This forms highly accurate statistical models of objects, processes and phenomena when there are vast quantities of data present or available.
In recent years, a variety of approaches for opening the deep and opaque models developed by AI training have appeared aiming to improve their interpretability and reliability. In the case of image recognition and LLMs, these concerns are largely a matter of ensuring an agreeable user experience. However, RL involves making decisions. In the context of games of pure entertainment, this relative lack of criticality of risks persists. However, when RL agents make decisions in contexts with real-world consequences, a notable degree of justified concern appears. An opaque black-box model simply cannot provide formal guarantees of safety or performance reliability. Rather, it is simply a model that efficiently provides a most likely answer, in some formal probabilistic sense. Low probability but catastrophic events could happen.
@@
In particular, crisp hard science-based models of system dynamics and foreseeable risks (e.g. of collision) are the standard, and a comprehensively reliable approach to endow RL systems with true guarantees of safe and stable operation. Said guarantees would significantly expand the set of domains and circumstances to which RL can be applied, and at the same time provide a transparent and interpretable means by which regulators can certify the safety of an RL system/agent. This will be done by embedding Model Predictive Control within the lower layer. Moreover, at the highest level of abstraction as to the agent’s operation, the decisions must be interpretable and associated with the mission and tasks associated with the AI. These should make logical sense to a user, and as far as the general agreement and sense of sequences of higher level operations. This will be done by embedding classical planning into the top layer.
In this paper we present a set of technical procedures from the methodological fields of process control and planning that incorporate structural assurances into deep neural networks in a systematic way. Along with addressing the aforementioned concerns of interpretability and reliability, these set of computational techniques significantly improve the quality and precision of the RL model and enable magnitudes of greater statistical power for any available datasets used to train the optimal decision model.
@@ 2 Hierarchical Reinforcement Learning
In this Section we provide the background for and the full formulation, as well as a description thereof, of a Bilevel Reinforcement Learning problem as it could be applied to the practical consideration above regarding elder care. To this end, we present a summary of the main principles of RL, its extension to continuous action spaces, and bilevel and generally hierarchical RL problems and solution algorithms.
@@ 2.1 Reinforcement Learning - Background
Formally, a standard RL problem (two prominent texts being sutton2018reinforcement ; bertsekas1996neuro ) is defined to consist of a set of states $\mathcal{S}$ and actions $\mathcal{A}$ to find a policy $\pi:\mathcal{X}\to\mathcal{M}(\mathcal{A})$ whereby $\mathcal{M}$ we mean the set of multinomial distributions over the set $\mathcal{A}$. The state evolves in time through an unknown transition probability tensor $\mathbf{P}$ with components $P(S_{t+1}=s^{\prime}|S_{t}=s,A_{t}=a)$. The reward $R(s,s^{\prime},a)$ defines some utility the agent receives upon that particular transition to $s^{\prime}$. In RL, techniques to learn the Value function $V^{\pi}(s)$ and Action-Value function $Q^{\pi}(s,a)$ are iterated together with establishing the policy $\pi$ in either performing the most rewarding action $a$ or randomly explore and learn more about the system.
@@
$\begin{split}&V^{\pi}(s(t))=\mathbb{E}_{a(t),s(t+1)}\left[R(s(t),a(t))+\gamma V^{\pi}(s(t+1))\right],\\ &Q^{\pi}(s(t),a(t))=\mathbb{E}_{s(t+1)}\left[R(s(t),a(t))+\gamma V^{\pi}(s(t+1))\right],\\ \end{split}$ (3)
@@ 2.3 Bilevel Reinforcement Learning Formulation
For Reinforcement Learning specifically kreidieh2019inter considers a boss-worker hierarchical RL with derivation of a cooperative policy gradient for the boss. The work zhu2022hierarchical defines navigation to avoid obstacles at a lower level, and goal-driven movement at the upper level, for a robotics RL engine. Image captioning with a language layer and a visual layer of RL is described in xu2019multi
@@
where $G$ is a functional that includes either $F$ and other local physical criteria, or tracking a solution that optimizes the top level functional.
@@ 3 Safe RL with Model Predictive Control
A significant challenge in applying RL to real-world problems is safety. Although approaches based on Constrained Markov Decision Processes (CMDPs) altman2021constrained ; achiam2017constrained allow minimization of the expected cumulative stage cost while following a set of safety constraints, they often require the exploration of unsafe trajectories during training. To avoid safety violations during training in particular (and to enhance learning efficiency in general), an emerging solution is the integration of domain knowledge or system constraints into the RL learning process garcia2015comprehensive .
The relevant domain for ensuring safe operation according to the laws of physics is Model Predictive Control (MPC). Indeed, the primary historical distinction between the two fields, model predictive control and reinforcement learning, developed under their separate domain native considerations, is the use of explicit background knowledge informed from known scientific principles, on the one hand, and (often very) large sample experiments querying a black box comprehensive roll-out simulation of the entire procedure. In practice, one is often in between the ideal scenario for MPC and R, that is, a partially complete physical understanding of some of the mechanics is available. As such, part of the system dynamics are known, and part have to be ascertained with data.
@@ 3.1 Model Predictive Control - Background
Formally, MPC consists of repeated solution of an optimal control problem (OCP) defined over some finite horizon as,
$\begin{array}[]{rl}x^{*},u^{*}\in&\arg\min\limits_{x(t)\in\mathbb{R}^{x},u(t)\in\mathcal{C}^{u}}\frac{1}{T}\sum\limits_{t=1}^{T}F(x(t),u(t))\\ \text{subject to }&x(t)=f(u(t),x(t-1))\\ &c^{I}(u(t),x(t))\leq 0,\,\forall t\\ &c^{E}(u(t),x(t))=0,\,\forall t\\ \end{array}$ (7)
where $f$ is a discretized differential equation (there are variants with implicit integrators as well, we simplify for presentation), and $c$ are constraints. There are control $u$ and state $x$ variables and $\mathcal{C}^{u}$ is some simple (e.g. bounds) constraint set
With MPC, one solves this problem, applies the first control, lets the system evolve, measures the state again, and solves a new OCP with the horizon shifted one step forward. Note two important distinctions from RL: 1) the lack of uncertain quantities or random variables in the model, and 2) the presence of functions defining constraints. These constraints can enforce safety, e.g., by indicating the geometry of the environment as far the presence of objects an agent can collide with.
@@
$\quad h\left(x_{t},u_{t}\right)\leq 0,\quad u_{t}\in\mathcal{A}\,,\quad$ (8c)
where $T$ is the prediction horizon, $F$ is the stage cost, $V_{T}$ is the terminal cost, $f$ is the dynamics and $h$ is the inequality constraint. Problem 8 produces a complete profile of control inputs $u^{\star}=\{u_{0}^{\star},\ldots,u_{N-1}^{\star}\}$ and corresponding state predictions $x^{\star}=\{x_{0}^{\star},\ldots,x_{N}^{\star}\}$. However, only the first element $u_{0}^{\star}$ of the input sequence $u^{\star}$ is applied to the system MPCbook . At the next physical sampling time, a new state $s$ is received, and problem 8 is solved again, producing a new $u^{\star}$ and a new $u_{0}^{\star}$. MPC (8) hence yields a policy
@@
with $u_{0}^{\star}$ solution of 8 for $s$ given. For $\gamma\approx 1$, policy (9) can provide a good approximation of the optimal policy $\pi^{\star}$ for an adequate choice of prediction horizon $T$, terminal cost $V_{N}$ and if the MPC model $f$ approximates the true dynamics sufficiently well.
@@ 3.2 Model Predictive Control and RL as Dynamic Programming
$c^{I}(u(t,\xi(1:t)),x(t,\xi(1:t)))\leq 0,\quad t=1,\dots,T,$
$c^{E}(u(t,\xi(1:t)),x(t,\xi(1:t)))=0,\quad t=1,\dots,T.$
@@ 3.3 Model Predictive Control Incorporation within RL
    $Q_{\theta^{\star}}(s,a)=Q^{\star}(s,a),\,\forall s\in\Omega$, for the inputs $a\in\mathcal{A}$ such that $\lvert V^{\star}(f_{\theta^{\star}}(s,a))\rvert<\infty$
@@
$\Omega$ =$=\left\{s\in\mathcal{S}\,\,\Big{|}\,\,\left|V^{\star}(x^{\star}_{t})\right|<\infty,\ \forall\,t\leq T\right\}\,,$ (17)
@@ 3.3.2 Real-time MPC
A key observation for real-time MPC is that the initial condition $s_{k}$ at a given discrete time step $k$ are typically close to their prediction formed at time $k-1$. More specifically, considering the predicted optimal state $x_{1}^{\star}(s_{k-1})\subset w^{\star}(s_{k-1})$ at time $k-1$ is typically a good approximation of $s_{k}$. This observation leads to the shifting strategy to provide an initial guess for NLP (19) at $s_{k}$ from the solution obtained at time $k-1$ for $s_{k-1}$. The shifting strategy consists in using the optimal state-input trajectories $x^{\star}_{0,\ldots,N}(s_{k-1})$ $u^{\star}_{0,\ldots,N-1}(s_{k-1})$ contained in $w^{\star}(s_{k})$ to form the state-input initial guesses:
@@
A key to ensuring that $x_{1}^{\star}(s_{k-1})\approx s_{k}$ is to make sure that the sampling time of the MPC scheme is fast enough to ensure that the prediction error cannot accumulate during the physical time interval between the discrete times $k-1$ and $k$. When this strategy is sufficiently effective, the SQP procedure using the initial guess (24) often fully converges in a very few iterations. An extreme version of SQP for real-time MPC is then to perform a single SQP iteration at each time sample $k$ of MPC, and rely on the strong quadratic contraction resulting from the update (23) from the initial guess (24)-(25) to keep the (not fully converged) primal-dual solutions close to their optimal values $w^{\star}(s),\lambda^{\star}(s),\mu^{\star}(s)$. This principle is at the core of the Real-Time Iteration (RTI). In addition to circumventing multiple SQP iterations, RTI seeks to minimize the time spent between obtaining a state measurement $s_{k}$ and delivering the corresponding approximation to the optimal solution $w^{\star}(s_{k})$. Indeed, delivering $w^{\star}(s_{k})$ requires performing two computationally expensive operations: forming QP (22), i.e. computing $H,\nabla_{w}\Phi(w),g(w,s_{k}),\nabla_{w}g(w,s),h(w),\nabla_{w}h(w)$, and solving it. A key observation behind the RTI is that it is possible to form QP (22) before receiving $s_{k}$. Hence the RTI performs the update of the MPC solution in two phases. The preparation phase consists in performing the shifting (24)-(25) and preparing QP (22), while the feedback phase consists in solving QP (22) once $s_{k}$ is received.

### v2

@@ 1 Introduction
In service of this task, the robot is able to lean to roll forward and back as well as turn in full rotation, and rotate its arm along three axes. Objects are picked up with grasping, or relevant task-oriented objects are magnetized. Objects in the environment define collisions to be avoided, which requires including the speed and position in the state space. The robot has a propulsion system, which is its forward control, and a braking mechanism. The manufacturing of this propulsion system involves some new features which are not system-identified to full certainty, and there is a small but noticeable nonlinear gyroscopic effect on its motion. This effect will have to be learned in the course of its operation.
This problem requires the incorporation of learning, as can be performed with standard RL procedures, together with tokenized interpretable actions and tasks for trustworthiness, together with physical mechatronic control of the robot’s actuators. This combination is famously handled in the present day with Vision-Language-Action models, e.g. ma2024survey . However, long term autonomous robotics for sensitive domains with potential critical risks, such as physical care, is still a long distance, in research and development output, from trustworthy operation at scale christoforou2020upcoming ; pandy2025ai . The fundamental problems associated with Deep RL for real world physical application is seminally summarized in dulac2019challenges . The problems observed in practice include a) poor out-of-distribution generalization, b) lack of formal guarantees of accuracy and performance, essential for trust in end users and regulators, and c) strict adherence to reasonable protocols of safety. These issues are fundamental: recall that RL is, ultimately, a statistical model, one that does not structurally incorporate the vast knowledge base of the natural sciences. As such, the mathematical representation of the model lacks the mechanistic intuition appropriate for understanding cause and effect kungurtsev2025cause and thus potential consequences of, e.g., hazardous actions.
Inspired by the recent flurry of activity of Reinforcement Learning Model Predictive Control (RLMPC) reiter2025synthesis , the present paper develops a methodology for incorporating domain knowledge in order to enhance the efficiency and reliability of autonomous deployment. Indeed, it has been observed in several applications that combining Reinforcement Learning with Model Predictive Control can achieve superior performance than both MPC and RL alone, see e.g. cai2023learning .
@@
We present a formalism that defines a bilevel optimization problem incorporating Reinforcement Learning (RL)-assisted Model Predictive Control in the bottom layer, and RL-assisted classical planning at the higher level. The overarching architecture of this proposed framework is depicted in Figure 1.
[…]
Figure 1: High-level methodology overview. A Scheduler (left) drives a Planner and MPC, ultimately controlling the physical Agent (right), with feedback loops returning state and reward information for respective updates.
@@
Human-aligned and safe autonomous robotics is a technology of significant interest to many researchers, companies, and governments, with population aging creating many potential labor shortages for assistance care and health for the elderly, the potential for automated manufacturing, and other reasons. As it is also a representative example of the conceptual components of such an operation, it will be a running example throughout this line of work, although the same principles apply to similar domains such as automated manufacturing, self-driving vehicles and other embodied AI. In this work we focus on a simplified stylistic model of long run autonomous robotic human care and focus on developing the foundations of the mathematical formulation for the appropriate optimization and data driven sequential decision making framework. We present a formal schema to define the autonomous decision procedures of an agent using multiple relevant disciplines associated with mathematical engineering.
@@
Designing and managing systems towards optimality or looser satisfaction criteria over time is a generic cross-disciplinary problem that has attracted attention from numerous communities. In the case wherein the model of the states and actions can be well represented with either a finite set or fairly low dimensional and bounded continuous domain while the system dynamics and the utility of different circumstances is unknown, Reinforcement Learning (RL) has risen to prominence as the preeminent methodology. RL, like Classification (e.g., image recognition), and Generative Learning (e.g., Large Language Models), has been a central component in the accelerating, groundbreaking developments of AI that have characterized the technological landscape for the past decade. Neural networks leverage the raw computational power of highly parallel GPU hardware permitting training of deep models with millions to billions of trainable parameters. This forms highly accurate statistical models of objects, processes and phenomena when there are vast quantities of data present or available.
In recent years, a variety of approaches for opening the deep and opaque models developed by AI training have appeared aiming to improve their interpretability and reliability. In the case of image recognition and LLMs, these concerns are largely a matter of ensuring an agreeable user experience. However, RL involves making decisions. In the context of games of pure entertainment, this relative lack of criticality of risks persists. However, when RL agents make decisions in contexts with real-world consequences, a notable degree of justified concern appears. An opaque black-box model simply cannot provide formal guarantees of safety or performance reliability. Rather, it is simply a model that efficiently provides a most likely answer, in some formal probabilistic sense. Low probability but catastrophic events could happen and the risk of such rare tail events is difficult to quantify – naturally given low sample counts for rare events, it is challenging to use purely empirical methods to predict and forecasts such risks and simulate contingency plans.
@@
In particular, crisp hard science-based models of system dynamics and foreseeable risks (e.g. of collision) are the standard, and a comprehensively reliable approach to endow RL systems with true guarantees of safe and stable operation. Said guarantees would significantly expand the set of domains and circumstances to which RL can be applied, and at the same time provide a transparent and interpretable means by which regulators can certify the safety of an RL system/agent. This will be done by embedding Model Predictive Control within the lower layer. Moreover, at the highest level of abstraction as to the agent’s operation, the decisions must be interpretable and associated with the mission and tasks associated with the AI. These should make logical and semantic sense to a user, and as far as the general agreement and sense of sequences of higher level operations. This will be done by embedding classical planning into the top layer.
In this paper we present a set of technical procedures from the methodological fields of process control and planning that incorporate structural assurances into deep neural networks in a systematic way. Along with addressing the aforementioned concerns of interpretability and reliability, these sets of computational techniques significantly improve the quality and precision of the RL model and enable magnitudes of greater statistical power for any available datasets used to train the optimal decision model.
In summary, the Contributions of this manuscript include:
  • A bilevel integrated data-driven optimization scheme for operationalizing high level interpretable token-based decisions together with low level physical control behavior,
  • A novel methodology for enriching Classical Scheduling and Planning with actions defining the operation of control problems and real-time streaming data adjusting the components Planning ontology,
  • A novel methodology of complete end-to-end Planning-Control differentiation through the use of fuzzification operators,
  • A systems engineering formalism that facilitates crisp guarantees of trustworthy interpretable actions and safe control behavior simultaneously with learning to facilitate adaptation to an uncertain and changing environment
  • Comprehensive validation through a standing example of Robotic Care throughout the manuscript together with a detailed numerical illustration.
@@ 2 Hierarchical Reinforcement Learning
In this Section we provide the background for and the full formulation, as well as a description thereof, of a Bilevel Reinforcement Learning problem as it could be applied to the practical consideration above regarding robotic human care. To this end, we present a summary of the main principles of RL, its extension to continuous action spaces, and bilevel and generally hierarchical RL problems and solution algorithms.
@@ 2.1 Reinforcement Learning - Background
Formally, a standard RL problem (two prominent texts being sutton2018reinforcement ; bertsekas1996neuro ) is defined to consist of a set of states $\mathcal{X}$ and a set of actions $\mathcal{A}$. The objective is to learn a policy $\pi:\mathcal{X}\to\mathcal{M}(\mathcal{A})$, where $\mathcal{M}(\mathcal{A})$ denotes the set of multinomial distributions over the action set $\mathcal{A}$. The state evolves in time through an unknown transition probability tensor $\mathbf{P}$ with components $P(S_{t+1}=s^{\prime}|S_{t}=s,A_{t}=a)$. The reward $R(s,s^{\prime},a)$ defines some utility the agent receives upon that particular transition to $s^{\prime}$. In RL, techniques to learn the Value function $V^{\pi}(s)$ and Action-Value function $Q^{\pi}(s,a)$ are iterated together with establishing the policy $\pi$ in either performing the most rewarding action $a$ or randomly explore and learn more about the system.
@@
[…]
@@ 2.3 Bilevel Reinforcement Learning Formulation
For Reinforcement Learning specifically kreidieh2019inter considers a boss-worker hierarchical RL with derivation of a cooperative policy gradient for the boss. The work zhu2022hierarchical defines navigation to avoid obstacles at a lower level, and goal-driven movement at the upper level, for a robotics RL engine. The captioning of the image with a language layer and a RL visual layer is described in xu2019multi
@@
where $G$ is a functional that includes either $F$ and other local physical criteria, or tracking a solution that optimizes the top level functional. This interaction between the slow-timescale discrete planner and the fast-timescale continuous controller is illustrated in Figure 2.
Upper Level RL (Discrete, Slow $t$) State $x(t)\in\mathcal{X}$: Discrete Context Action $a(t)\in\mathcal{A}$: Feed, Clean, … Policy $\pi:\mathcal{X}\to\mathcal{A}$ $\min\mathbb{E}\left[\sum\gamma^{t}F(\dots)\right]$ Lower Level Control (Continuous, Fast $\tau$) State $z(\tau)$: Location, Velocity Control $v(\tau)$: Motor Inputs Policy $\zeta:z(\tau)\to v(\tau)$ $\min\mathbb{E}\left[\sum\beta^{\tau}G(\dots)\right]$ Environment & Rewards Evaluate Costs $F$ and $G$ + Reward Accomplished- Penalty CollisionAction a(t)a(t)Control $v(\tau)$Discrete State x(t+1)x(t+1)Continuous State $z(\tau+h)$
Figure 2: Hierarchical Bilevel RL Architecture: Visualizing the decoupling of timescales. The upper level policy assigns discrete abstract actions (e.g., feeding a human) to the lower level. The lower level evaluates continuous state feedback at a high frequency to generate physical control outputs. The resulting environmental trajectory determines the reward signals evaluated by the cost functions $F$ and $G$.
@@ 3 Safe RL with Model Predictive Control
A significant challenge in applying RL to real-world problems is safety. Although approaches based on Constrained Markov Decision Processes (CMDPs) altman2021constrained ; achiam2017constrained allow minimization of the expected cumulative stage cost while following a set of safety constraints, they often require the exploration of unsafe trajectories during training. To avoid safety violations during training in particular (and to enhance learning efficiency in general), an emerging solution is the integration of domain knowledge or system constraints into the RL learning process garcia2015comprehensive . This approach has been shown to outperform data-driven RL methods with soft penalties discouraging constraint violation, e.g. ray2019benchmarking .
The relevant domain for ensuring safe operation according to the laws of physics is Model Predictive Control (MPC). Indeed, the primary historical distinction between the two fields, model predictive control and reinforcement learning, developed under their separate domain native considerations, is the use of explicit background knowledge informed from known scientific principles, on the one hand, and (often very) large sample experiments querying a black box comprehensive roll-out simulation of the entire procedure. In practice, one is often in between the ideal scenario for MPC and RL, that is, a partially complete physical understanding of some of the mechanics is available. As such, part of the system dynamics are known, and part have to be ascertained with data.
@@ 3.1 Model Predictive Control - Background
Formally, MPC consists of repeated solution of an optimal control problem (OCP) defined over some finite horizon.
With MPC, one solves the discretized nonlinear optimization problem defining an approximation of the near future system behavior, applies the first control, lets the system evolve, measures the state again, and solves a new OCP with the horizon shifted one step forward. Note two important distinctions from RL: 1) the lack of uncertain quantities or random variables in the model, and 2) the presence of functions defining constraints. These constraints can enforce safety, e.g., by indicating the geometry of the environment as far the presence of objects an agent can collide with.
@@
$\quad h\left(x_{t},u_{t}\right)\leq 0,\quad u_{t}\in\mathcal{A}\,$$\quad$ (7c)
where $T$ is the prediction horizon, $F$ is the stage cost, $V_{T}$ is the terminal cost, $f$ is the dynamics and $h$ is the inequality constraint. Problem 7 produces a complete profile of control inputs $u^{\star}=\{u_{0}^{\star},\ldots,u_{N-1}^{\star}\}$ and corresponding state predictions $x^{\star}=\{x_{0}^{\star},\ldots,x_{N}^{\star}\}$. However, only the first element $u_{0}^{\star}$ of the input sequence $u^{\star}$ is applied to the system MPCbook . At the next physical sampling time, a new state $s$ is received, and problem 7 is solved again, producing a new $u^{\star}$ and a new $u_{0}^{\star}$. MPC (7) hence yields a policy
@@
with $u_{0}^{\star}$ solution of 7 for $s$ given. For $\gamma\approx 1$, policy (8) can provide a good approximation of the optimal policy $\pi^{\star}$ for an adequate choice of prediction horizon $T$, terminal cost $V_{N}$ and if the MPC model $f$ approximates the true dynamics sufficiently well.
@@ 3.2 Model Predictive Control and RL as Dynamic Programming
$h(u(t,\xi(1:t)),x(t,\xi(1:t)))\leq 0,\quad t=1,\dots,T,$
@@ 3.3 Model Predictive Control Incorporation within RL
    $Q_{\theta^{\star}}(s,a)=Q^{\star}(s,a),\,\forall s\in\Omega$, for the inputs $a\in\mathcal{A}$ such that $\lvert V^{\star}(f_{\theta^{\star}}(s,a))\rvert$$\infty$
@@
$\Omega$ $:=\left\{s\in\mathcal{S}\,\,\Big|\,\,\left|V^{\star}(x^{\star}_{t})\right|<\infty,\ \forall\,t\leq T\right\}\,$, (16)
@@ 3.3.2 Real-time MPC
A key observation for real-time MPC is that the initial condition $s_{k}$ at a given discrete time step $k$ are typically close to their prediction formed at time k1k-1. More specifically, considering the predicted optimal state $x_{1}^{\star}(s_{k-1})\subset w^{\star}(s_{k-1})$ at time k1k-1 is typically a good approximation of $s_{k}$. This observation leads to the shifting strategy to provide an initial guess for NLP (18) at $s_{k}$ from the solution obtained at time k1k-1 for $s_{k-1}$. The shifting strategy consists in using the optimal state-input trajectories $x^{\star}_{0,\ldots,N}(s_{k-1})$ $u^{\star}_{0,\ldots,N-1}(s_{k-1})$ contained in $w^{\star}(s_{k})$ to form the state-input initial guesses:
@@
A key to ensuring that $x_{1}^{\star}(s_{k-1})\approx s_{k}$ is to make sure that the sampling time of the MPC scheme is fast enough to ensure that the prediction error cannot accumulate during the physical time interval between the discrete times k1k-1 and $k$. When this strategy is sufficiently effective, the SQP procedure using the initial guess (23) often fully converges in a very few iterations. An extreme version of SQP for real-time MPC is then to perform a single SQP iteration at each time sample $k$ of MPC, and rely on the strong quadratic contraction resulting from the update (22) from the initial guess (23)-(24) to keep the (not fully converged) primal-dual solutions close to their optimal values $w^{\star}(s),\lambda^{\star}(s),\mu^{\star}(s)$. This principle is at the core of the Real-Time Iteration (RTI). In addition to circumventing multiple SQP iterations, RTI seeks to minimize the time spent between obtaining a state measurement $s_{k}$ and delivering the corresponding approximation to the optimal solution $w^{\star}(s_{k})$. Indeed, delivering $w^{\star}(s_{k})$ requires performing two computationally expensive operations: forming QP (21), i.e. computing $H,\nabla_{w}\Phi(w),g(w,s_{k}),\nabla_{w}g(w,s),h(w),\nabla_{w}h(w)$, and solving it. A key observation behind the RTI is that it is possible to form QP (21) before receiving $s_{k}$. Hence the RTI performs the update of the MPC solution in two phases. The preparation phase consists in performing the shifting (23)-(24) and preparing QP (21), while the feedback phase consists in solving QP (21) once $s_{k}$ is received.